# DeFi Arbitrage in Hedged Liquidity Tokens

### v2

@@ 5 Discussion
To overcome the aforementioned issue in which the $2\sqrt{P}$ pricing at other CPMMs (such as at Uniswap pools) limits the liquidity available under variable minting costs, instead we can consider variable fee rates instead. That is, consider a CPMM in which the fees charged are provided by the strictly decreasing mapping $\gamma:\mathbb{R}_{++}\to\mathbb{R}_{++}$ of the outstanding liquidity tokens $L>0$. Now, instead of updating the market price $2\sqrt{P}$, the CPMM has an updated risk-neutral valuation $V_{0}(P,L)=2\hat{\gamma}(L)\sqrt{P}/\hat{\gamma}^{*}$ per liquidity token. We recommend that $\gamma_{0}:=\lim_{L\searrow 0}\gamma(L)$ is set sufficiently small (e.g., $30bps$ as was used in all Uniswap v2 pools) so that arbitrageurs would act in approximation to the schema introduced herein and $\lim_{L\nearrow\infty}\gamma(L)=0$ so that any reasonable price level can be supported. For example, $\gamma(L)=\gamma_{0}\exp(-\alpha L)$ or $\gamma(L)=\gamma_{0}(1+\alpha L)^{-1}$ for parameter $\alpha>0$ provides a control over the fee dependence on liquidity. Following the logic of Section 4.1, an implied volatility can be given that now has explicit dependence on the level of liquidity $L$ in the CPMM pool. The primary drawback to this construction is that the LP cannot know the realized fee rate for their investment when making the purchase; instead she will need to continuously review this investment to determine if it remains advantageous.

### v3

@@ 5 Discussion
To overcome the aforementioned issue in which the $2\sqrt{P}$ pricing at other CPMMs (such as at Uniswap pools) limits the liquidity available under variable minting costs, we can consider variable fee rates instead. That is, consider a CPMM in which the fees charged are provided by the strictly decreasing mapping $\gamma:\mathbb{R}_{++}\to\mathbb{R}_{++}$ of the outstanding liquidity tokens $L>0$. Now, instead of updating the market price $2\sqrt{P}$, the CPMM has an updated risk-neutral valuation $V_{0}(P,L)=2\hat{\gamma}(L)\sqrt{P}/\hat{\gamma}^{*}$ per liquidity token as per Theorem 3.2. Recall that $\hat{\gamma}^{*}$, as defined in (3.4), is constant for fixed $r,\sigma$ and is independent of the liquidity $L$. With this construction, the ratio $\hat{\gamma}(L)/\hat{\gamma}^{*}$ provides a delineation of whether LPing is a good investment or not, assuming all future fees remain constant at $\gamma(L)$. As opposed to the constant fee setting, as investors may have different beliefs about volatility $\sigma$, they each estimate $\hat{\gamma}^{*}$ differently leading to equilibrium liquidity $L^{*}$; in this way, if an investor believe that $\hat{\gamma}(L^{*})<\hat{\gamma}^{*}$ then they will not invest while an investor that believes that $\hat{\gamma}(L^{*})>\hat{\gamma}^{*}$ will choose to increase the CPMM liquidity. We recommend that $\gamma_{0}:=\lim_{L\searrow 0}\gamma(L)$ is set sufficiently small (e.g., $30bps$ as was used in all Uniswap v2 pools) so as to encourage LPing and $\lim_{L\nearrow\infty}\gamma(L)=0$ so that any reasonable price level can be supported. For example, $\gamma(L)=\gamma_{0}\exp(-\alpha L)$ or $\gamma(L)=\gamma_{0}(1+\alpha L)^{-1}$ for parameter $\alpha>0$ provides a control over the fee dependence on liquidity. Following the logic of Section 4.1, an implied volatility can be given that now has explicit dependence on the level of liquidity $L$ in the CPMM pool. The primary drawback to this construction is that the LP cannot know the realized fee rate for their investment when making the purchase; instead she will need to continuously review this investment to determine if it remains advantageous.